OpenStudy (anonymous):

Inverse functions: I'm having trouble figuring out what this means (in terms of which way to begin solving): f(x) = 10x + Sqrt(x+101) what is (f^(-1))'(0) =? that's (d/dx)(inverse)(0).. I don't get it.. do you find the inverse of f, then the derivative and then plug in zero? If so can you please show me how you solve this? Thank you!

OpenStudy (anonymous):

ok lets make sure you are not confused in to thinking you can actually find the inverse of this you cannot, so don't try it

OpenStudy (anonymous):

what you are going to use is that \[\frac{d}{dx}[f^{-1}(0)]=\frac{1}{f'(f^{-1}(0))}\]

OpenStudy (anonymous):

ok, i was trying.. thanks for telling me this

OpenStudy (anonymous):

as you see from the above, you do not need the derivative of the inverse, you need only know two things: the inverse of 0, and the derivative of the original function \(f\) evaluated at that number

OpenStudy (anonymous):

so first lets find the inverse of 0 by inspection \[f^{-1}(0)\] is the number you would plug in to \(10x+\sqrt{x+101}\) to get \(0\) what do you think that number is? (hint, don't do algebra, just think)

OpenStudy (anonymous):

still thinking?

OpenStudy (anonymous):

yes, because the only way i can think of doing this is with algebra.. and im forcing my self to not htink that way.. this is bad haha

OpenStudy (anonymous):

im goign to just say -101..

OpenStudy (anonymous):

or 0

OpenStudy (anonymous):

oh heck no 0 doesn't work because if \(x=0\) you get \(10\times 0+\sqrt{101}=\sqrt{101}\) think again

OpenStudy (anonymous):

10x^2 +x+101.. and just solve for x?

OpenStudy (anonymous):

lol.. ok too much pressure.. hold on

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

think that is all what would make this beast zero?

OpenStudy (anonymous):

oh crap.. it looks much more complicated.. like (10x + sqrt(x+101))^2

OpenStudy (anonymous):

it is not supposed to be hard you have \(\sqrt{x+101}\) what would you put for \(x\) to make a very simple square inside there?

OpenStudy (anonymous):

stop doing algebra

OpenStudy (anonymous):

i can tell you are doing algebra way too much work suppose i asked you what would make \(2x+\sqrt{x+5}=0\) would that be easier?

OpenStudy (anonymous):

kind of.. i feel like you would take 1 away from 100 (or 5).. like x+100 +1..?

OpenStudy (anonymous):

i don't mean to torture you i will tell you if you like, but then you will say "oh doh"

OpenStudy (anonymous):

x+4+1

OpenStudy (anonymous):

yes, you would take one away from 101! so what would you make \(x\) to get \(\sqrt{100}\) in there?

OpenStudy (anonymous):

i have \(\sqrt{x+101}\) and i wish it was \(\sqrt{100}\) what does that make \(x\)??

OpenStudy (anonymous):

haha.. let me take some time.. i want to be sure im right.. joking, x=10

OpenStudy (anonymous):

no not ten

OpenStudy (anonymous):

so it's 10sqrt(x+1)

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

ok i am done torturing you if i have \(\sqrt{x+101}\) and i wish it was \(\sqrt{100}\) i would set \(100=x+101\) and solve for \(x\) in one step

OpenStudy (anonymous):

oh, i was answering the sqrt(100).. sorry.. is the 10sqrt(x+1 right?

OpenStudy (anonymous):

solve \[100=x+101\]

OpenStudy (anonymous):

oh.. i see.. x=-1

OpenStudy (anonymous):

WHEW!!

OpenStudy (anonymous):

haha.. ok i guess im not a lost cause!

OpenStudy (anonymous):

now lets check that it is right \[f(x)=10x+\sqrt{x+101}\] \[f(-1)=10\times (-1)+\sqrt{-1+101}=-10+10=0\] we got it

OpenStudy (anonymous):

no not at all but you should look for the more or less obvious answer here ok now that we know \[f^{-1}(0)=-1\] we can continue next thing we need is \(f'(x)\)

OpenStudy (anonymous):

ohhhh... ok i see.. so with these problems.. you can usually see the relationship between whats under the radical and whats outside of it.. so like 10^2 is 100.. and so i would just find that x=-1.. without the amount of energy it took this time.. nice

OpenStudy (anonymous):

if you have not seen this kind of problem before, i can imagine that you would think you have to do something difficult, but usually they make the answer obvious (more or less) now \[f(x)=10x+\sqrt{x+101}\] what is \(f'(x)\) ?

OpenStudy (anonymous):

10+(1/2)((x+101)^(-1/2))

OpenStudy (anonymous):

10+1/2sqrt(x+101)

OpenStudy (anonymous):

ok good, but i don't like it because of the exponential notation you should memorize the fact that the derivative of \(\sqrt{x}\) is \(\frac{1}{2\sqrt{x}}\)without resorting to the power rule it is a very common function, and the derivative never changes so once you have it memorized, like knowing that \(8\times 9=72\) you can just move on

OpenStudy (anonymous):

so we got \[f'(x)=10+\frac{1}{2\sqrt{x+101}}\] and we also know \(f^{-1}(0)=-1\) next job is to find \[f'(-1)\]

OpenStudy (anonymous):

yeah i started seeing the relationship the more i did it.. i think i know the answer now.. f'(-1) =10..?

OpenStudy (anonymous):

because 10+(1/2*sqrt(100)) = 10+1/20.. sorry i can do math.. 200+1/20

OpenStudy (anonymous):

ok good \[f'(x)=10+\frac{1}{2\sqrt{x+101}}\] \[f'(-1)=10+\frac{1}{2\sqrt{-1+101}}=10+\frac{1}{20}=\frac{201}{20}\]

OpenStudy (anonymous):

and final job is to flip it

OpenStudy (anonymous):

since it's 1/f'(f^-1).. so it's 20/201?

OpenStudy (anonymous):

aka take the reciprocal yes

OpenStudy (anonymous):

now lets quickly recap, because we spent a lot of time, so lets get exactly what we wanted, and what we needed okay?

OpenStudy (anonymous):

when you wrote the equation (different from what I was given) how did you know to make the denominator equal to the inverse(0)? i want to say its because f(x)' or in this case f(inv)prime is related to dx?

OpenStudy (anonymous):

okay yeah

OpenStudy (anonymous):

here we go we have \[f(x)=10x+\sqrt{x+101}\] and you want \[(f^{-1})'(0)\] we are going to use \[(f^{-1})')(0)=\frac{1}{f'(f^{-1}(0))}\] for this we need \[f^{-1}(0)\] which, after some work, we find \(f^{-1}(0)=-1\) then we need \[f'(x)=10+\frac{1}{2\sqrt{x+101}}\] and \[f'(-1)=\frac{201}{20}\] putting it all together we get \[(f^{-1})'(0)=\frac{1}{f'(f^{-1}(0))}=\frac{1}{f'(-1)}=\frac{1}{\frac{201}{20}}=\frac{20}{201}\]

OpenStudy (anonymous):

Got it.. i just redid it from the top with different numbers and i got them right (online)! Wohoo

OpenStudy (anonymous):

Thank you so much! i have an exam coming up so this has been really helpful!

OpenStudy (anonymous):

afwan

OpenStudy (anonymous):

haha.. not arab but kind gesture.. :)

OpenStudy (anonymous):

but you knew what it meant, right?

OpenStudy (anonymous):

yeah.. i hang around arabs.. shukran is thank you in both languages (mine is urdu).. and i'm not sure how to say you're welcome in urdu.. need to learn that =/

OpenStudy (anonymous):

hmmm if your mother is like mine, she is very disappointed that you do not know her native language i surely don't my arabic is limited to please, thank you, and food

OpenStudy (anonymous):

the problem with urdu is it's comprised of like a bunch of different languages.. like arabic, hindi, farsi and all of the related ones.. my parents speak different dialects so it was just easier to speak english at home.. but i was brought up learning urdu (till age 4 or 5).. but i really want to learn farsi/arabic.. mom knows farsi since she was born near the boarder

OpenStudy (anonymous):

arabic is awesome.. you should learn it by watching tv in arabic!

OpenStudy (anonymous):

my father is american, so only english at home i thought about it, but it is hard, new alphabet, rules etc i'll stick with the food good luck on your exam, i am sure you will do well

OpenStudy (anonymous):

food is always good.. and thanks! i'll prob be posting more problems soon haha